# Certified Defense on the Fairness of Graph Neural Networks

### v2

@@ Abstract
Graph Neural Networks (GNNs) have emerged as a prominent graph learning model in various graph-based tasks over the years. Nevertheless, due to the vulnerabilities of GNNs, it has been empirically proved that malicious attackers could easily corrupt the fairness level of their predictions by adding perturbations to the input graph data. In this paper, we take crucial steps to study a novel problem of certifiable defense on the fairness level of GNNs. Specifically, we propose a principled framework named ELEGANT and present a detailed theoretical certification analysis for the fairness of GNNs. ELEGANT takes any GNNs as its backbone, and the fairness level of such a backbone is theoretically impossible to be corrupted under certain perturbation budgets for attackers. Notably, ELEGANT does not have any assumption over the GNN structure or parameters, and does not require re-training the GNNs to realize certification. Hence it can serve as a plug-and-play framework for any optimized GNNs ready to be deployed. We verify the satisfactory effectiveness of ELEGANT in practice through extensive experiments on real-world datasets across different backbones of GNNs, where ELEGANT is also demonstrated to be beneficial for GNN debiasing. Open-source code can be found at https://github.com/yushundong/ELEGANT.
1 Introduction
@@
Graph Neural Networks (GNNs) have emerged among the most popular models to handle learning tasks on graphs [32, 46] and made remarkable achievements in various domains [16, 35]. Nevertheless, as GNNs are increasingly deployed in real-world decision-making scenarios, there has been an increasing societal concern on the fairness of GNN predictions. A primary reason is that most traditional GNNs do not consider fairness, and thus could exhibit bias against certain demographic subgroups. Here the demographic subgroups are usually divided by certain sensitive attributes, such as gender and race. To prevent GNNs from biased predictions, multiple recent studies have proposed fairness-aware GNNs [1, 9, 28] such that potential bias could be mitigated.
@@
Unfortunately, despite existing efforts towards fair GNNs, it remains difficult to prevent the corruption of their fairness level due to their common vulnerability of lacking adversarial robustness. In fact, malicious attackers can easily corrupt the fairness level of GNNs by perturbing the node attributes (i.e., changing the values of node attributes) and/or the graph structure (i.e., adding and deleting edges) [20], which could lead to serious consequences in the test phase [9, 20]. For example, GNNs have been leveraged to perform bail decision-making on the graph of defendants, where an edge between two defendants represents high profile similarity [1]. Yet, by simply injecting adversarial links in the graph data, attackers can make GNNs deliver advantaged predictions for a subgroup (e.g., individuals with a certain nationality) while damaging the interest of others [20]. Hence achieving defense over the fairness of GNNs is crucial for the purpose of safe deployment.
@@
It is worth noting that despite the abundant empirical defense strategies for GNNs [52, 14, 22, 25, 48], they are always subsequently defeated by novel attacking techniques [43, 6], and the defense over the fairness of GNNs also faces the same problem. Therefore, an ideal way is to achieve certifiable defense on fairness (i.e., certified fairness defense). A few recent works aim to certify the fairness for traditional deep learning models [30, 29, 23, 38, 5, 42]. Nevertheless, most of them require specially designed training strategies [30, 23, 42] and thus cannot be directly applied to optimized GNNs ready to be deployed. More importantly, they mostly rely on assumptions on the optimization results [30, 23, 5, 42] or data distributions [29, 38] over a continuous input space. Hence they can hardly be generalized to GNNs due to the binary nature of the input graph topology. Several other works propose certifiable GNN defense approaches to achieve theoretical guarantee [47, 3, 4, 21, 55, 56]. However, they mainly focus on securing the GNN prediction for a certain individual node to ensure model utility, ignoring the fairness defense over the entire population. Therefore, despite the significance, the study in this field still remains in its infancy.
@@
It is worth noting that achieving certifiable defense on the fairness of GNNs is a daunting task due to the following key challenges: (1) Generality: different types of GNNs could be designed and optimized for different real-world applications [53]. Correspondingly, our first challenge is to design a plug-and-play framework that can achieve certified defense on fairness for any optimized GNN models that are ready to be deployed. (2) Vulnerability: a plethora of existing studies have empirically verified that most GNNs are sensitive to input data perturbations [52, 54, 49]. In other words, small input perturbations may cause significant changes in the GNN output. Hence our second challenge is to properly mitigate the common vulnerabilities of GNNs without changing its structure or re-training. (3) Multi-Modality: the input data of GNNs naturally bears multiple modalities. For example, there are node attributes and graph topology in the widely studied attributed networks. In practice, both data modalities may be perturbed by malicious attackers. Therefore, our third challenge is to achieve certified defenses of fairness on both data modalities for GNNs.
@@
As an early attempt to address the aforementioned challenges, in this paper, we propose a principled framework named ELEGANT (cEtifiabLE GNNs over the fAirNess of PredicTions). Specifically, we focus on the widely studied task of node classification and formulate a novel research problem of Certifying GNN Classifiers on Fairness. To handle the first challenge, we propose to develop ELEGANT on top of an optimized GNN model without any assumptions over its structure or parameters. Hence ELEGANT is able to serve as a plug-and-play framework for any optimized GNN model ready to be deployed. To handle the second challenge, we propose to leverage randomized smoothing [47, 8] to defend against malicious attacks, where most GNNs can then be more robust over the prediction fairness level. To handle the third challenge, we propose two different strategies working in a concurrent manner, such that certified defense against the attacks on both the node attributes (i.e., add and subtract attribute values) and graph topology (i.e., flip the existence of edges) can be realized. Finally, we evaluate the effectiveness of ELEGANT on multiple real-world network datasets. In summary, our contributions are three-fold: (1) Problem Formulation. We formulate and make an initial investigation on a novel research problem of Certifying GNN Classifiers on Fairness. (2) Algorithm Design. We propose a framework ELEGANT to achieve certified fairness defense against attacks on both node attributes and graph structure without relying on assumptions about any specific GNNs. (3) Experimental Evaluation. We perform comprehensive experiments on real-world datasets to verify the effectiveness of ELEGANT.
2 Problem Definition
@@
Preliminaries. Let $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ be an undirected attributed network, where $\mathcal{V}=\{v_{1},...,v_{n}\}$ is the set of $n$ nodes; $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. Let $\bm{A}\in\{0,1\}^{n\times n}$ and $\bm{X}\in\mathbb{R}^{n\times d}$ be the adjacency matrix and attribute matrix of $\mathcal{G}$, respectively. Assume each node in $\mathcal{G}$ represents an individual, and sensitive attribute $s$ divides the population into different demographic subgroups. We follow a widely studied setting [1, 9] to assume the sensitive attribute is binary, i.e., $s\in\{0,1\}$. We use $s_{i}$ to denote the value of the sensitive attribute for node $v_{i}$. In node classification tasks, we use $\mathcal{V}_{\text{trn}}$ and $\mathcal{V}_{\text{tst}}$ ($\mathcal{V}_{\text{trn}},\mathcal{V}_{\text{tst}}\in\mathcal{V}$) to represent the training and test node set, respectively. We denote the GNN node classifier as $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$. $f_{\bm{\theta}}$ takes $\bm{A}$ and $\bm{X}$ as input, and outputs $\hat{\bm{Y}}$ as the predictions for the nodes in $\mathcal{G}$. Each row in $\hat{\bm{Y}}$ is a one-hot vector flagging the predicted class. We use $f_{\bm{\theta}^{*}}$ to denote the GNN with optimal parameter $\bm{\theta}^{*}$.
@@
Threat Model. We focus on the attacking scenario of model evasion, i.e., the attack happens in the test phase. In particular, we assume that the victim model under attack is an optimized GNN node classifier $f_{\bm{\theta}^{*}}$. We follow a widely adopted setting [3, 55, 37, 39] to assume that a subset of nodes $\mathcal{V}_{\text{vul}}\in\mathcal{V}_{\text{tst}}$ are vulnerable to attacks. Specifically, attackers may perturb their links (i.e., flip the edge existence) to other nodes and/or their node attributes (i.e., change their attribute values). We denote the perturbations on adjacency matrix as $\bm{A}\oplus\bm{\Delta}_{\bm{A}}$. Here $\oplus$ denotes the element-wise XOR operator; $\bm{\Delta}_{\bm{A}}\in\{0,1\}^{n\times n}$ is the matrix representing the perturbations made by the attacker, where 1 only appears in rows and columns associated with the vulnerable nodes while 0 appears elsewhere. Correspondingly, in $\bm{\Delta}_{\bm{A}}$, 1 entries represent edges that attackers intend to flip, while 0 entries are associated with edges that are not attacked. Similarly, we denote the perturbations on node attribute matrix as $\bm{X}+\bm{\Delta}_{\bm{X}}$, where $\bm{\Delta}_{\bm{X}}\in\mathbb{R}^{n\times n}$ is the matrix representing the perturbations made by the attacker. Usually, if the total magnitude of perturbations is within certain budgets (i.e., $\|\bm{\Delta}_{\bm{A}}\|_{0}\leq\epsilon_{\bm{A}}$ for $\bm{A}$ and $\|\bm{\Delta}_{\bm{X}}\|_{2}\leq\epsilon_{\bm{X}}$ for $\bm{X}$), the perturbations are regarded as unnoticeable. The goal of an attacker is to add unnoticeable perturbations to nodes in $\mathcal{V}_{\text{vul}}$, such that the GNN predictions for nodes in $\mathcal{V}_{\text{tst}}$ based on the perturbed graph exhibit as much bias as possible. In addition, we assume that the attacker has access to any information about the victim GNN (i.e., a white-box setting). This is the worst case in practice, which makes it even more challenging to achieve defense.
@@ 3 Methodology
3.1 Bias Indicator Function
@@
(Bias Indicator Function) Given adjacency matrix $\bm{A}$ and node attribute matrix $\bm{X}$, a test node set $\mathcal{V}_{\text{tst}}$, a threshold $\eta$ for the exhibited bias, and an optimized GNN model $f_{\bm{\theta}^{*}}$, the bias indicator function is defined as $g(f_{\bm{\theta}^{*}},\bm{A},\bm{X},\eta,\mathcal{V}_{\text{tst}})=\mathbbm{1}\left(\pi(f_{\bm{\theta}^{*}}(\bm{A},\bm{X}),\mathcal{V}_{\text{tst}})<\eta\right)$, where $\mathbbm{1}(\cdot)$ takes an event as input and outputs 1 if the event happens (otherwise 0); $\pi(\cdot,\cdot)$ denotes any bias metric for GNN predictions (taken as its first parameter) over a set of nodes (taken as its second parameter). Traditional bias metrics include $\Delta_{\text{SP}}$ [9, 13] and $\Delta_{\text{EO}}$ [9, 19].
@@
3.2 Certified Fairness Defense over Node Attributes
@@
We denote $\bm{\Gamma}_{\bm{X}}=\gamma_{\bm{X}}(\bm{\omega}_{\bm{X}},\mathcal{V}_{\text{vul}})$ and $g(\bm{A},\bm{X}+\bm{\Gamma}_{\bm{X}})=g(f_{\bm{\theta}^{*}},\bm{A},\bm{X}+\gamma_{\bm{X}}(\bm{\omega}_{\bm{X}},\mathcal{V}_{\text{vul}}),\eta,\mathcal{V}_{\text{tst}})$ below for simplicity. We are then able to derive the theoretical certification for the defense on fairness with the defined $\tilde{g}_{\bm{X}}$ in Definition 2. Below we present a certification for the defense of fairness directly building upon the certification given in [8].
@@
(Certified Fairness Defense for Node Attributes [8]) Denote the probability for $g(\bm{A},\bm{X}+\bm{\Gamma}_{\bm{X}})$ to return class $c$ ($c\in\{0,1\}$) as $P(c)$. Then $\tilde{g}_{\bm{X}}(\bm{A},\bm{X})$ will provably return $\mathrm{argmax}_{c\in\{0,1\}}P(c)$ for any perturbations (over the attributes of vulnerable nodes) within an $l_{2}$ radius $\tilde{\epsilon_{\bm{X}}}=\frac{\sigma}{2}\left(\Phi^{-1}(\max_{c\in\{0,1\}}P(c))-\Phi^{-1}(\min_{c\in\{0,1\}}P(c))\right)$, where $\Phi^{-1}(\cdot)$ is the inverse of the standard Gaussian cumulative distribution function.
Correspondingly, for an $\eta$ that enables $\max_{c\in\{0,1\}}P(c)=1$, it is then safe to say that no matter what perturbations $\bm{\Delta}_{\bm{X}}$ are made on vulnerable nodes, as long as $\|\bm{\Delta}_{\bm{X}}\|_{2}\leq\tilde{\epsilon}_{\bm{X}}$, the constructed $\tilde{g}_{\bm{X}}$ will provably not yield predictions for $\mathcal{V}_{\text{tst}}$ with a level of bias exceeding $\eta$. Nevertheless, it is worth noting that, in GNNs, perturbations may also be made on the structure of the vulnerable nodes, i.e., adding and/or deleting edges between these vulnerable nodes and any nodes in the graph. Hence it is also necessary to achieve certified defense against such structural attacks. Here we propose to also smooth the constructed $\tilde{g}_{\bm{X}}$ over the graph structure (of the vulnerable nodes) for the purpose of certified fairness defense on the graph structure. However, the adjacency matrix describing the graph structure is naturally binary, and thus should be smoothed in a different way.
3.3 Certified Fairness Defense over Node Attributes and Graph Structure

### v3

@@ Abstract.
Graph Neural Networks (GNNs) have emerged as a prominent graph learning model in various graph-based tasks over the years. Nevertheless, due to the vulnerabilities of GNNs, it has been empirically shown that malicious attackers could easily corrupt the fairness level of their predictions by adding perturbations to the input graph data. In this paper, we take crucial steps to study a novel problem of certifiable defense on the fairness level of GNNs. Specifically, we propose a principled framework named ELEGANT and present a detailed theoretical certification analysis for the fairness of GNNs. ELEGANT takes any GNN as its backbone, and the fairness level of such a backbone is theoretically impossible to be corrupted under certain perturbation budgets for attackers. Notably, ELEGANT does not make any assumptions over the GNN structure or parameters, and does not require re-training the GNNs to realize certification. Hence it can serve as a plug-and-play framework for any optimized GNNs ready to be deployed. We verify the satisfactory effectiveness of ELEGANT in practice through extensive experiments on real-world datasets across different backbones of GNNs and parameter settings.
Graph Neural Networks; Algorithmic Fairness
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780323isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Machine learning
1. Introduction
@@
Graph Neural Networks (GNNs) have emerged among the most popular models to handle learning tasks on graphs (DBLP:conf/iclr/KipfW17; velivckovic2017graph) and made remarkable achievements in various domains (feng2022twibot; li2022graph). Nevertheless, as GNNs are increasingly deployed in real-world decision-making scenarios, there has been an increasing societal concern on the fairness of GNN predictions. A primary reason is that most traditional GNNs do not consider fairness, and thus could exhibit bias against certain demographic subgroups. Here the demographic subgroups are usually divided by certain sensitive attributes, such as gender and race. To prevent GNNs from biased predictions, multiple recent studies have proposed fairness-aware GNNs (agarwal2021towards; dai2021say; kangwww2022; dong2023reliant) such that potential bias could be mitigated.
@@
Unfortunately, despite existing efforts towards fair GNNs, it remains difficult to prevent the corruption of their fairness level due to their common vulnerability of lacking adversarial robustness. In fact, malicious attackers can easily corrupt the fairness level of GNNs by perturbing the node attributes (i.e., changing the values of node attributes) and/or the graph structure (i.e., adding and deleting edges) (hussain2022adversarial), which could lead to serious consequences in the test phase (dai2021say; hussain2022adversarial). For example, GNNs have been leveraged to perform bail decision-making on the graph of defendants, where an edge between two defendants represents high profile similarity (agarwal2021towards). Yet, by simply injecting adversarial links in the graph data, attackers can make GNNs deliver advantaged predictions for a subgroup (e.g., individuals with a certain nationality) while damaging the interest of others (hussain2022adversarial). Hence achieving defense over the fairness of GNNs is crucial for the purpose of safe deployment.
@@
It is worth noting that despite the abundant empirical defense strategies for GNNs (zhang2020gnnguard; entezari2020all; jin2019latent; jin2020graph; wu2019adversarial), they are always subsequently defeated by novel attacking techniques (schuchardt2020collective; carlini2017adversarial), and the defense over the fairness of GNNs also faces the same problem. Therefore, an ideal way is to achieve certifiable defense on fairness (i.e., certified fairness defense). A few recent works aim to certify the fairness for traditional deep learning models (khedr2022certifair; kang2022certifying; jin2022input; mangold2022differential; borca2022provable; ruoss2020learning). Nevertheless, most of them require specially designed training strategies (khedr2022certifair; jin2022input; ruoss2020learning) and thus cannot be directly applied to optimized GNNs ready to be deployed. More importantly, they mostly rely on assumptions on the optimization results (khedr2022certifair; jin2022input; borca2022provable; ruoss2020learning) or data distributions (kang2022certifying; mangold2022differential) over a continuous input space. Hence they can hardly be generalized to GNNs due to the binary nature of the input graph topology. Several other works propose certifiable GNN defense approaches to achieve theoretical guarantee (wang2021certified; bojchevski2019certifiable; bojchevski2020efficient; jin2020certified; zugner2019certifiable; zugner2020certifiable). However, they mainly focus on securing the GNN prediction for a certain individual node to ensure model utility, while how to achieve fairness defense over the entire population is ignored. Therefore, despite the significance, the study in this field still remains in its infancy.
@@
In fact, achieving certifiable defense on the fairness of GNNs is a daunting task due to the following key challenges: (1) Generality: different types of GNNs could be designed and optimized for different real-world applications (zhou2020graph). Correspondingly, our first challenge is to design a plug-and-play framework that can achieve certified defense on fairness for any optimized GNN models that are ready to be deployed. (2) Vulnerability: a plethora of existing studies have empirically verified that most GNNs are sensitive to input data perturbations (zhang2020gnnguard; zugner2020adversarial; xu2019topology). In other words, small input perturbations may cause significant changes in the GNN output. Hence our second challenge is to properly mitigate the common vulnerabilities of GNNs without changing its structure or re-training. (3) Multi-Modality: the input data of GNNs naturally bears multiple modalities. For example, there are node attributes and graph topology in the widely studied attributed networks. In practice, both data modalities may be perturbed by malicious attackers. Therefore, our third challenge is to achieve certified defenses of fairness on both data modalities for GNNs.
@@
As an early attempt to address the aforementioned challenges, in this paper, we propose a principled framework named ELEGANT (cEtifiabLE GNNs over the fAirNess of PredicTions). Specifically, we focus on the widely studied task of node classification and formulate a novel research problem of Certifying GNN Classifiers on Fairness. To handle the first challenge, we propose to develop ELEGANT on top of an optimized GNN model without any assumptions over its structure or parameters. Hence ELEGANT is able to serve as a plug-and-play framework for any optimized GNN model ready to be deployed. To handle the second challenge, we propose to leverage randomized smoothing (wang2021certified; cohen2019certified) to defend against malicious attacks, where most GNNs can then be more robust over the prediction fairness level. To handle the third challenge, we propose two different strategies working in a concurrent manner, such that certified defense against the attacks on both the node attributes (i.e., add and subtract attribute values) and graph topology (i.e., flip the existence of edges) can be realized. Finally, we evaluate the effectiveness of ELEGANT on multiple real-world network datasets. In summary, our contributions are three-fold:
  • Problem Formulation. We first formulate and perform an initial investigation on the novel research problem of Certifying GNN Classifiers on Fairness.
  • Algorithm Design. We propose a framework ELEGANT to achieve certified fairness defense against attacks on both node attributes and graph structure without relying on assumptions about any specific GNNs.
  • Experimental Evaluation. We perform comprehensive experiments on real-world datasets under different GNN backbones to verify the effectiveness of ELEGANT.
2. Problem Definition
@@
Preliminaries. Let $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ be an undirected attributed network, where $\mathcal{V}=\{v_{1},...,v_{n}\}$ is the set of $n$ nodes; $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. Let $\bm{A}\in\{0,1\}^{n\times n}$ and $\bm{X}\in\mathbb{R}^{n\times d}$ be the adjacency matrix and attribute matrix of $\mathcal{G}$, respectively. Assume each node in $\mathcal{G}$ represents an individual, and sensitive attribute $s$ divides the population into different demographic subgroups. We follow a widely studied setting (agarwal2021towards; dai2021say) to assume the sensitive attribute is binary, i.e., $s\in\{0,1\}$. We use $s_{i}$ to denote the value of the sensitive attribute for node $v_{i}$. In node classification tasks, we use $\mathcal{V}_{\text{trn}}$ and $\mathcal{V}_{\text{tst}}$ ($\mathcal{V}_{\text{trn}},\mathcal{V}_{\text{tst}}\in\mathcal{V}$) to represent the training and test node set, respectively. We denote the GNN node classifier as $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$. $f_{\bm{\theta}}$ takes $\bm{A}$ and $\bm{X}$ as input, and outputs $\hat{\bm{Y}}$ as the predictions for the nodes in $\mathcal{G}$. Each row in $\hat{\bm{Y}}$ is a one-hot vector flagging the predicted class. Furthermore, we use $f_{\bm{\theta}^{*}}$ to denote a GNN equipped with the optimal parameter $\bm{\theta}^{*}$.
@@
Threat Model. We focus on the attacking scenario of model evasion, i.e., the attack happens in the test phase. In particular, we assume that the victim model under attack is an optimized GNN node classifier $f_{\bm{\theta}^{*}}$. We follow a widely adopted setting (bojchevski2019certifiable; zugner2019certifiable; ma2020towards; mu2021hard) to assume that a subset of nodes $\mathcal{V}_{\text{vul}}\in\mathcal{V}_{\text{tst}}$ are vulnerable to attacks. Specifically, attackers may perturb their links (i.e., flip the edge existence) to other nodes and/or their node attributes (i.e., change their attribute values). We denote the perturbations on adjacency matrix as $\bm{A}\oplus\bm{\Delta}_{\bm{A}}$. Here $\oplus$ denotes the element-wise XOR operator; $\bm{\Delta}_{\bm{A}}\in\{0,1\}^{n\times n}$ is the matrix representing the perturbations made by the attacker, where 1 only appears in rows and columns associated with the vulnerable nodes while 0 appears elsewhere. Correspondingly, in $\bm{\Delta}_{\bm{A}}$, 1 entries represent edges that attackers intend to flip, while 0 entries are associated with edges that are not attacked. Similarly, we denote the perturbations on node attribute matrix as $\bm{X}+\bm{\Delta}_{\bm{X}}$, where $\bm{\Delta}_{\bm{X}}\in\mathbb{R}^{n\times d}$ is the matrix representing the perturbations made by the attacker. Usually, if the total magnitude of perturbations is within certain budgets (i.e., $\|\bm{\Delta}_{\bm{A}}\|_{0}\leq\epsilon_{\bm{A}}$ for $\bm{A}$ and $\|\bm{\Delta}_{\bm{X}}\|_{F}\leq\epsilon_{\bm{X}}$ for $\bm{X}$), the perturbations are regarded as unnoticeable. The goal of an attacker is to add unnoticeable perturbations to nodes in $\mathcal{V}_{\text{vul}}$, such that the GNN predictions for nodes in $\mathcal{V}_{\text{tst}}$ based on the perturbed graph exhibit as much bias as possible. In addition, we assume that the attacker has access to any information about the victim GNN (i.e., a white-box setting). We note that such assumption aligns with the worst case in practice, which makes it even more challenging to achieve defense.
@@ 3. Methodology
3.1. Bias Indicator Function
@@
(Bias Indicator Function) Given adjacency matrix $\bm{A}$ and node attribute matrix $\bm{X}$, a test node set $\mathcal{V}_{\text{tst}}$, a threshold $\eta$ for the exhibited bias, and an optimized GNN model $f_{\bm{\theta}^{*}}$, the bias indicator function is defined as $g(f_{\bm{\theta}^{*}},\bm{A},\bm{X},\eta,\mathcal{V}_{\text{tst}})=\mathbf{1}\left(\pi(f_{\bm{\theta}^{*}}(\bm{A},\bm{X}),\mathcal{V}_{\text{tst}})<\eta\right)$, where $\mathbf{1}(\cdot)$ takes an event as input and outputs 1 if the event happens (otherwise 0); $\pi(\cdot,\cdot)$ denotes any bias metric for GNN predictions (taken as its first parameter) over a set of nodes (taken as its second parameter). Traditional metrics measuring the exhibited algorithmic bias include $\Delta_{\text{SP}}$ (dai2021say; DBLP:conf/innovations/DworkHPRZ12) and $\Delta_{\text{EO}}$ (dai2021say; DBLP:conf/nips/HardtPNS16).
@@
3.2. Certified Defense over Node Attributes
@@
In the discussion below, we denote $\bm{\Gamma}_{\bm{X}}=\gamma_{\bm{X}}(\bm{\omega}_{\bm{X}},\mathcal{V}_{\text{vul}})$ and $g(\bm{A},\bm{X}+\bm{\Gamma}_{\bm{X}})=g(f_{\bm{\theta}^{*}},\bm{A},\bm{X}+\gamma_{\bm{X}}(\bm{\omega}_{\bm{X}},\mathcal{V}_{\text{vul}}),\eta,\mathcal{V}_{\text{tst}})$ for simplicity. Based on Definition 2, we are now able to derive the theoretical certification for the defense on fairness with the defined $\tilde{g}_{\bm{X}}$ in Definition˜2. Below we present a certification for the defense of fairness directly building upon the certification given in (cohen2019certified).
@@
(Certified Fairness Defense for Node Attributes (cohen2019certified)) Denote the probability for $g(\bm{A},\bm{X}+\bm{\Gamma}_{\bm{X}})$ to return class $c$ ($c\in\{0,1\}$) as $P(c)$. Then $\tilde{g}_{\bm{X}}(\bm{A},\bm{X})$ will provably return $\mathrm{argmax}_{c\in\{0,1\}}P(c)$ for any perturbations (over the attributes of vulnerable nodes) within an $l_{2}$ radius $\tilde{\epsilon_{\bm{X}}}=\frac{\sigma}{2}\left(\Phi^{-1}(\max_{c\in\{0,1\}}P(c))-\Phi^{-1}(\min_{c\in\{0,1\}}P(c))\right)$, where $\Phi^{-1}(\cdot)$ is the inverse of the standard Gaussian cumulative distribution function.
Correspondingly, for an $\eta$ that enables $\max_{c\in\{0,1\}}P(c)=1$, it is then safe to say that no matter what perturbations $\bm{\Delta}_{\bm{X}}$ are made on vulnerable nodes, as long as $\|\bm{\Delta}_{\bm{X}}\|_{F}\leq\tilde{\epsilon}_{\bm{X}}$, the constructed $\tilde{g}_{\bm{X}}$ will provably not yield predictions for $\mathcal{V}_{\text{tst}}$ with a level of bias exceeding $\eta$. Nevertheless, it is worth noting that, in GNNs, perturbations may also be made on the structure of the vulnerable nodes, i.e., adding and/or deleting edges between these vulnerable nodes and any nodes in the graph. Hence it is also necessary to achieve certified defense against such structural attacks. Here we propose to also smooth the constructed $\tilde{g}_{\bm{X}}$ over the graph structure (of the vulnerable nodes) for the purpose of certified fairness defense on the graph structure. However, the adjacency matrix describing the graph structure is naturally binary, and thus should be smoothed in a different way. We elaborate on the joint certification over node attributes and graph structure in the next subsection.
3.3. Certified Defense over Node Attributes and Graph Structure